# Less is More : A Closer Look at Multi-Modal Few-Shot Learning

### 10 Jan 2024 v1

@@ 5.3.1 Ablation study
MLP $86\mathrm{K}$ $\mathbf{75.27}$ $\mathbf{83.89}$ $\mathbf{84.92}$ $\mathbf{89.10}$
@@
Table 3 exhibits the ablation study on four datasets under the 5-way 1-shot setting. By combining the visual backbone with the pre-trained LM even with the fixed Prompts, the accuracy can be improved obviously. Then the utilization of the learnable prompts instead the fixed prompts (SimpleFSL) also bring the obvious performance improvements. Furthermore, we adopt the self-ensemble and self-distillation module (SimpleFSL++) further improved the accuracy on all four datasets. On the whole, the proposed modules used in SimpleFSL and SimpleFSL++ contribute to the performance. Additionally, the ablation study also demonstrates that utlizing explicitly the pre-trained LM is significant for the FSL.
5.3.2 Prompts analysis
As discussed above, we introduce the learnable prompts instead of fiexed prompts, and ablation study demonstrates its effectiveness for FSL. The learnable prompts we used in SimpleFSL and SimpleFSL++ are all dataset-aware, which means the learned vector will shared with all classes in a dataset. Inspired by recent progresses in prompt Learning [60, 59], we test the class-aware prompts, which means the learnable prompts is related the classes and conditional on the visual prototypes. Further, considering the few-shot learning consists of many few-shot tasks, we proposed the task-aware prompts, which means the learnable prompt is related the few-shot tasks and conditional on the visual features of all samples in a few-shot learning task. Table 4 summarize the 5-way 1/5-shot accuracy comparisons about the three variants of learnable prompts. And in Table 4, the fine granularity of modeling from top to bottom decreases sequentially. Unfortunately, the deployments of task-aware and class-aware prompts both do not lead to better performance. These phenomena also consist with the conclusion in previous research [60], that the more fine-grained modeling of prompts may not boost performances in the downstream tasks. And these may caused by the overfitting because the more fine-grained modeling, the more parameters to learning, which is particularly challenging in low-data scenarios. Consequently, we adopt the dataset-aware prompts in our framework and how to design a more suitable prompts for FSL deserves further investigation. The more experimental details and settings about the designs of class-aware prompt and task-aware prompt can be found in the Appendix.
5.3.3 Adaptor analysis
The adaptor in multi-modal few-shot learning plays a significant role [51, 6] which transfers the textual representation into the visual representation space. As shown in Table 5, we empirically observed that the choose of different some impact to the classification performance. The liner adaptor are wildly used in previous works [51, 6]. The 2-layer MLP with the bottleneck structure are used in our framework, the hidden dimension of which is much smaller than others. Inspired by the NLP adaptor [19], we also evaluate the combination of the bottleneck structure and a liner layer with the residual connection [16]. The experimental results demonstrate that our MLP w/ bottleneck has slight advantage, which may own to its less parameters. The experimental details about the three adaptors can be found in the Appendix.
@@ 5.3.3 Adaptor analysis
Add $75.27$ $\mathbf{83.89}$ $\mathbf{84.92}$ $\mathbf{89.07}$
@@
Attention [51] $\mathbf{75.28}$ $83.75$ $84.39$ $\mathbf{89.07}$
5.3.4 Fusion mechanism
Both in SimpleFSL and SimpleFSL++, we adpot the simple add operation to fuse the visual representation and textual representation for the Multi-modal Feature Fusion mechanism. We also re-implement Simple++ with different fusion mechanism, including the Concatenation and Attention [51], and experimental results are summarized in Table 6. We observe that the choice of Add or Attention has a little impact on the performance. For simplicity, we adopt the add operation in our framework. The experimental details about the Concatenation and Attention can be found in the Appendix.
@@ 5.3.5 Hyper-parameters analysis
As shown in the Ablation study, the self-Distillation module brings the additional boost to the performance. And we make a more detailed evaluation of the impact on the performance in SimpleFSL++. Figure 3 shows the results of 5-way 1/5-shot learning on the miniImageNet and CIFAR-FS. We observe that the utilization of self-distillation module with an appropriate weight improves the performance on both two datasets, while too-large a value of weight may not bring the boost.
@@ 6 Conclusion
In this paper, we emphasize the generalization capability of the pre-trained language model in the few-shot learning taks, which has been underestimated in previous works. Consequently, we propose a simple and effective framework for few-shot learning tasks, and we explicitly exploit the generalization capability of the pre-trained LM with the learnable prompt. Further, we apply the self-ensemble and self-Distillation to further bring the additional performance boost. Our extensive experiments conducted across four few-shot datasets demonstrate that our simple framework achieves promising results. especially in the 1-shot learning task. How to design the more suitable prompts for FSL deserves further investigation in the future.

### 24 Mar 2024 v2

@@ 5.3.3 Adaptor analysis
MLP $86\mathrm{K}$ $\mathbf{75.59}$ $\mathbf{83.89}$ $\mathbf{85.09}$ $\mathbf{89.10}$
@@
The adaptor in semantic-based few-shot learning plays a significant role [58, 6] which transfers the textual representation into the visual representation space. As illustrated in Fig.5, we empirically observe that different adaptors bear some influence on the classification performance. The linear adaptor is frequently used in previous works [58, 6]. The 2-layer MLP with a bottleneck structure are used in our framework, with a small number of hidden dimension. Inspired by the NLP adaptors [20], we also evaluate the combination of the bottleneck structure and a linear layer with the residual connection [18]. The experimental results demonstrate that our MLP w/ bottleneck has slight advantage, which may own to its less parameters. The experimental details about the three adaptors can be found in the Appendix.
5.3.4 Fusion mechanism
@@
SP-CLIP [6] $72.41$ $83.23$ $82.18$ $88.24$
Add $75.59$ $\mathbf{83.89}$ $\mathbf{85.09}$ $\mathbf{89.10}$
@@
Attention [58] $\mathbf{75.61}$ $83.75$ $84.39$ $\mathbf{89.10}$
In both SimpleFSL and SimpleFSL++, we adopt a straightforward addition operation to fuse the visual representation and semantic representation for multi-modal feature fusion. We also re-implement Simple++ with different fusion mechanism, including the Concatenation and Attention [58], and experimental results are summarized in Table 6. Here we use the SP-CLIP [6] as the baseline , which design a complex fusion mechanisms to insert semantic representations into the feature extractor. The results in Fig.6 reveal negligible differences in performance between employing Addition or Attention. This suggests that the simple fusion operation is benefit for the utilization of pre-trained LMs, and the obtained semantic feature can directly assist to classify without passing the complex architecture which may hurt the generalization capability of semantic features. For simplicity, we adopt the add operation in our framework. The experimental details about the Concatenation and Attention are described in the Appendix.
@@ 5.3.5 Hyper-parameters analysis
Figure 3: Evaluation of different weighting factor $\lambda$ on miniImageNet and CIFAR-FS.
@@
The ablation study highlights that the self-ensemble and self-distillation module both confer an additional performance enhancement. Then, we delve deeper into the impact of these modules on the performance in SimpleFSL++. As delineated previously, $\lambda$ servers as the weighting factor to control the weighting of two classifiers in Eq.7, and $\alpha$ denotes the weighting factor of the distillation loss in Eq.9. Fig.3 and Fig.4 summarize the performance with varied $\lambda$ and $\alpha$ under 5-way 1/5-shot learning on the miniImageNet and CIFAR-FS, respectively. We simultaneously observe that the utilization of self-ensemble and distillation module with an appropriate weight ameliorates performance on these datasets. While an excessively high weighting value may not yield further benefits. Specially, an observation of Fig.3 reveals that the ideal $\lambda$ setting for 1-shot learning is lower compared to that in 5-shot learning. This discrepancy can potentially be also attributed to the fact that 5-shot learning offers a more stonger visual supervision signal in comparison to 1-shot learning, thereby diminishing the relative importance of semantic guidance, as discussed in Sec 5.2.
Figure 5: t-SNE visualizations of the representation of different methods.
5.3.6 Visualization
In Fig.5, we present the t-SNE [37] visualizations of the last-layer representation prior to classifier-3 in SimpleFSL++, compared to a sole visual backbone without semantic guidance and SP-CLIP [6]. We visualize the support set in a randomly sampled 5-way 200-shot task from $\mathbf{D_{novel}}$ in miniImageNet and CIFAR-FSL, with distinct classes denoted by different colors. All plots in Fig.5 collectively demonstrate that ours SimpleFSL++ achieves more distinct class separability, notwithstanding its straightforward architecture. In comparison, the more intricate fusion mechanism of SP-CLIP does not significantly surpass the discriminability provided by a visual backbone without semantic guidance.
@@ 6 Conclusion
In this paper, we focus on the few-shot image classification task, and emphasize the generalization capability of the pre-trained language model in the few-shot learning, which is usually underestimated in previous works. To harness this capability, we propose a straightforward and efficacious framework for few-shot learning tasks, instead of designing the intricate and complex architectures. And we directly add the visual feature with the textual feature with adaptable learnable prompts, allowing for flexible accommodation to various datasets. Further, we apply the self-ensemble and self-Distillation to bring the additional performance boost. Our extensive experiments conducted across four few-shot datasets demonstrate that our proposed framework consistently delivers promising results, with particularly notable performance in the 1-shot learning task. The exploration of prompt design, tailored to optimize few-shot learning deserves further investigation in the future.
Supplementary Material
Appendix 0.A Ablation Study of pre-trained LM
In this section, we aim to explore the impact that various pre-trained Language Models (LMs) exert on our framework’s performance, as have done in previous Semantic-based few-shot learning works [58, 34, 60, 6]. As delineated in the methodology section, we utilize pre-trained LMs as the semantic feature extractors, and the compared LMs include: GloVe [41], ALBERT [27], RoBERTa [35] and CLIP [44]. For a fair comparison, we deploy these LMs with fixed prompts (e.g., a photo of a cat) here in our proposed SimpleFSL framework. As reported in Tab.S1, we first observe that our proposed framework consistently bring the significantly performance improvement with various LMs, compared to the purely visual backbone without using LMs. This outcome demonstrably affirms the efficacy of our framework. Moreover, we observe that incorporating CLIP leads to superior performance in most scenarios compared to the other LMs. This observation is in harmony with the findings of recent studies [60, 6], which may be attributable to that CLIP’s features particularly well-suited for aligning semantic and visual representations. Therefore, we opt CLIP as our default semantic feature extractor in our implementations.
Table S1: Performance comparison with different language models in SimpleFSL on miniImageNet and CIFAR-FS.
Language Model miniImageNet CIFAR-FS
1-shot 5-shot 1-shot 5-shot
w.o. LM 65.16 81.22 71.99 85.98
GloVe [41] $67.89$ $81.73$ $80.37$ $88.57$
ALBERT [27] $67.31$ $81.50$ $80.20$ $88.08$
RoBERTa [35] $70.11$ $82.26$ $\mathbf{81.26}$ $88.72$
CLIP [44] $\mathbf{72.78}$ $\mathbf{83.34}$ $81.12$ $\mathbf{88.86}$
Appendix 0.B Additional Visualization
In the section of Visualization, we have presented the t-SNE visualizations of representations of different methods, sampled from the miniImageNet and CIFAR-FS datasets. And we provide more visualization examples from the tieredImageNet and FC100 datasets. Similarly, we visualize the support set in a randomly sampled 5-way 200-shot task from $\mathbf{D_{novel}}$ in tieredImageNet and FC100, with each class represented by a unique color. Notably, we observe that ours SimpleFSL++ continues to yield impressive performance in these two datasets, presenting more distinct features in comparison to other baselines. These visualizations reinforce our earlier observation that the more intricate fusion mechanism utilized by SP-CLIP [6] fails to offer a significant advantage in discriminability over a visual backbone without semantic guidance.
Figure S1: t-SNE visualizations of the representation of different methods.
Appendix 0.C Few-Shot Baselines with LM
In this section, we describe the details of two simple Few-Shot Baselines with pre-trained language models (LM) we designed, as have shown in Figure 1 of Section 1, and we name them with zero-shot and zero-shot+LP, respectively. Overall, we adopt the CLIP-like [44] pre-training methodology, which sololy utilize the base dataset for training, as depicted in Figure S2, and we do not utilize any labels from the novel dataset following the zero-shot setting [43].
Zero-shot. The illustration of zero-shot is shown in Figure S2(a). For the visual branch, we feed the images into the visual backbone to extract the visual feature. For the textual branch, we feed the prompts into the textual backbone to extract the textual feature. Following previous works [44, 6], the input prompt we chose is "A photo of a [classname]", and we employ the textual encoder of CLIP as our textual backbone. We also utilize an adaptor subsequent to the textual backbone to transform textual representations into the visual representation space. During the training stage, we aim to maximize the cosine similarity between the corresponding visual features and textual features [44]. We only update the parameters of the visual backbone and the adaptor, while keeping the textual backbone frozen. During inference, we follow the zero-shot setting [44], and directly predict the samples in the novel dataset without leveraging any label.
Zero-shot+LP. Building upon the zero-shot baseline, the zero-shot+LP is depicted in Figure S2(b). We adopt the learnable prompts instead of the pre-defined fixed prompts, while training and inference processes remain consistent with those of the zero-shot baseline. The parameters of the learnable prompts are updated during the training stage, and the textual backbone is still frozen. We adopt the dataset-aware prompt here as introduced in Section 4.2.
Figure S2: Illustration of the used Few-Shot Baselines with LM in Section 1, concluding the zero-shot and zero-shot+LP.
Appendix 0.D Details of Prompt Analysis
As previously discussed in Section 5.3.2, our framework incorporates the dataset-aware learnable prompts. Additionally, we also evaluate the performances of two alternative prompt settings: class-aware and task-aware prompts. Then we introduce the detailed configuration of these two variants of learnable prompt.
Class-aware Prompt. Inspired by recent advancements in prompt Learning [68, 67], we attempt to utilize the learnable class-aware prompts, which are dynamically tailored to specific classes and conditional on the visual class prototypes. Formally, we obtain the class-aware prompt $\mathbf{v}_{class}$ by utilizing both the dataset-aware prompt $\mathbf{v}$ and the visual prototype $\boldsymbol{p_{i}^{0}}$ as follows:
$\mathbf{v}_{class}=\mathbf{v}+\pi(\boldsymbol{p_{i}^{0}})$ (1)
where $\pi(\ )$ denotes the class-aware Net, which is designed to generate the conditional prompt vectors for each prototype. And we employ a two-layer MLP as the class-aware Net here. Then, the obtained dataset-aware prompt $\mathbf{v}$ is fed into the textual backbone for subsequent processes.
Task-aware Prompt. Considering the particularity of meta-training strategy, which contains numerous few-shot learning (FSL) tasks, with each FSL task containing distinct samples and classes, we also investigate the performance of the task-aware prompt, denoted with $\mathbf{v}_{task}$. Analogous to the class-aware prompt, we obtain $\mathbf{v}_{task}$ by utilizing the dataset-aware prompt $\mathbf{v}$ and all visual prototype $\boldsymbol{p_{i}^{0}}$ within an few-shot learning task:
$\mathbf{v}_{task}=\mathbf{v}+\pi(\Sigma_{i}^{K}\boldsymbol{p_{i}^{0}})$ (2)
Similar to the class-aware Net, $\pi(\ )$ here denotes the task-aware Net, which aims to generate conditional prompt vectors for each FSL task, for which we also utilize a two-layer MLP.
Appendix 0.E Details of Adaptor Analysis
Figure S3: Illustration of three adaptors utilized in the paper
In this section, we provide details about the compared adaptors, as discussed in Section 5.3.3, containing the linear adaptor, MLP with the bottleneck, and the NLP adaptor [20]. And we visualize their structures in Figure S3, respectively.
Linear adaptor. The linear adaptor is prevalently utilized in previous works [58, 6]. The transformed textual semantic feature $z$ can be obtained by:
$z=\boldsymbol{W_{0}}\ g(y^{text}_{\mathbf{v}})+\boldsymbol{b_{0}}$ (3)
where $\boldsymbol{W_{0}}$ and $\boldsymbol{b_{0}}$ are the learnable parameters of the adaptor module. And the resulting $z\in\mathbb{R}^{d_{v}}$ has the same dimension with the visual feature $f(\boldsymbol{x_{j}})$.
MLP with the bottleneck. We adopt a two-layer Multi-Layer Perceptron (MLP) with bottleneck structure as the adaptor in our framework. This transformation process is expressed as:
$z=\boldsymbol{W_{2}}\ \sigma(\boldsymbol{W_{1}}\ g(y^{text}_{\mathbf{v}})+\boldsymbol{b_{1}})+\boldsymbol{b_{2}}$ (4)
where $\boldsymbol{W_{1}}$, $\boldsymbol{W_{2}}$, $\boldsymbol{b_{1}}$, and $\boldsymbol{b_{2}}$ are the learnable parameters. The intermediate hidden dimension we set is $\frac{d_{v}}{4}$.
NLP adaptor. Inspired by the NLP adaptor [20], we also evaluate the combination of the bottleneck structure and a linear layer with the residual connection [18]. An illustration of this structure can be found in Figure S3(c), and formally, we have:
$z^{\prime}$ $=\boldsymbol{W_{0}}\ g(y^{text}_{\mathbf{v}})+\boldsymbol{b_{0}}$ (5)
$z^{\prime\prime}$ $=\boldsymbol{W_{2}}\ \sigma(\boldsymbol{W_{1}}\ z^{\prime})+\boldsymbol{b_{1}})+\boldsymbol{b_{2}}$ (6)
$z$ $=z^{\prime}+z^{\prime\prime}$ (7)
where equation (5) and (6) denote the linear structure and MLP, respectively. And the resulting $z\in\mathbb{R}^{d_{v}}$ in equation (7) has the same dimension with the visual feature $f(\boldsymbol{x})$. The intermediate hidden dimension of the MLP we set is $\frac{d_{v}}{4}$, and Equation (7) depicts the residual connection [18].
Appendix 0.F Details of Fusion Mechanism
In this section, we delve into the specifics of the differing Fusion mechanisms discussed in Section 5.3.4, particularly focusing on Concatenation and Attention mechanisms.
Concatenation. Recalling that we have obtained the visual feature $f(\boldsymbol{x})$ and the transformed textual semantic feature $z$, the Concatenation fusion mechanism is articulated as:
$z_{f}=\boldsymbol{W_{f}}\ (f(\boldsymbol{x})||z)$ (8)
where $\boldsymbol{W_{f}}$ denotes the learnable parameters in the Fusion mechanism, and ||||| | means the Concatenation operation. The generated fusion feature $z_{f}$ shares the same dimension as the visual feature $f(\boldsymbol{x})$.
Attention. We aim to learn the fusion weight automaticly with an attention mechanism. Formally, we have:
$\alpha$ $=\mathbf{Att}(f(\boldsymbol{x})||z)\in(0,1)$ (9)
$z_{f}$ $=\alpha z+(1-\alpha)f(\boldsymbol{x})$ (10)
where $\alpha$ represents the learned fusion weight, and $\mathbf{Att}(\ )$ refers to the attention mechanism which consist of a two-layer MLP with a sigmoid activation.
As discussed in Section 5.3.4, we empirically observe that the choice of Addition or Attention marginally affects the performance. For the sake of simplicity, our framework adopts the additive operation.